OpenStudy (anonymous):

Calculate the derivative of the function s(x) = 3x − 9 2x + 4 3

OpenStudy (anonymous):

\[thats (\frac{ 3x-9 }{ 2x+4 })^{3}\]

OpenStudy (anonymous):

I got \[(\frac{ 3x-9 }{ 2x+4 })^{2}(\frac{ 9x-27 }{ (2x+4)^{2} })\] but thats incorrect, i cant find my mistake

zepdrix (zepdrix):

\[\Large \left[\left(\frac{ 3x-9 }{ 2x+4 }\right)^{3}\right]'\quad=\quad 3\left(\frac{ 3x-9 }{ 2x+4 }\right)^{2}\color{royalblue}{\left(\frac{3x-9}{2x+4}\right)'}\]Hmm let's see what the quotient rule gives us. I think you may have made a little boo boo somewhere.

zepdrix (zepdrix):

\[\Large \color{royalblue}{=\quad\frac{(3x-9)'(2x+4)-(3x-9)(2x+4)'}{(2x+4)^2}}\] \[\Large \color{royalblue}{=\quad\frac{(3)(2x+4)-(3x-9)(2)}{(2x+4)^2}}\]

OpenStudy (anonymous):

@zepdrix oh so i combine the two? So the answer is what typed last?

OpenStudy (anonymous):

i thought i used quotient rule on the right part and power rule for the left, then wrote both for the answer

zepdrix (zepdrix):

This is the answer before quotient rule has been applied to the blue part:\[\Large 3\left(\frac{ 3x-9 }{ 2x+4 }\right)^{2}\color{royalblue}{\left(\frac{3x-9}{2x+4}\right)'}\]Then the blue part is expanded below, plug the blue back into the blue :) It's kinda lengthy so I was trying to stay organized by not rewriting everything. Yes the answer is both parts. Your answer was very close, you shouldn't have gotten a `9x-27` in your numerator though. I think it works out to `30` instead.

OpenStudy (anonymous):

@zepdrix 30 for the numerator doesn't work either, thats still incorrect somehow

zepdrix (zepdrix):

Hmm ok sec :x

zepdrix (zepdrix):

Hmm :(

OpenStudy (anonymous):

@zepdrix was 3 multiplied into 30 already??